### v1

# Development and Benchmarking of JANGOFETT: A Novel Geant4-Operated Fission Event Tracking Tool

Abstract
Experiments measuring fission observables encounter false coincidences arising from timing overlap of separate fission product decays. Simulations of both fission observables and particles in detector systems exist, but have not yet been combined to produce accurate event-by-event outputs in a time-dependent manner. Geant4 is a powerful simulation tool for nuclear physics studies, but it does not handle multiple initial particles in a single simulation instance, nor does it feature high fidelity fission sampling. JANGOFETT: A Novel Geant4-Operated Fission Event Tracking Tool ) has been developed to address this challenge. The tool utilizes simulated fission data from an external program in conjunction with Geant4, which has been modified to produce a single timeline of events over an entire simulated experiment. The physical accuracy of the simulated overlapping energy depositions within detectors has been verified via simulation of fission products from the spontaneous fission of 252Cf.
keywords:
Geant4 , Fission , Nuclear Physics , Simulation , Radiation
journal: Nuclear Instruments and Methods in Physics Research Section A Accelerators, Spectrometers, Detectors, and Associated Equipment
[…]
organization=Oregon State University, Department of Nuclear Science and Engineeringequalequalfootnotetext: These authors contributed equally to this work. correspcorrespfootnotetext: Corresponding author. E-mail address: chemeya@oregonstate.edu.
1 Introduction
There is significant interest in simulating fission and its subsequent detection in experimental apparatus (Randrup and Vogt, ). Numerous experimental detection systems have been developed to correlate fission data at accelerator facilities around the world. Extensive simulations typically precede experiments with complex radiation detection systems to evaluate their potential and inform experiment sensitivity. This is often regarded as a necessary benchmark before any experiments occur, and it provides a valuable baseline for analysis code development.
There exist excellent software packages to facilitate Monte Carlo simulations assessing nuclear reactions and particle interactions within detectors such as MCNP (Los Alamos Scientific Laboratory. Group X-6, ) and Geant4 (Agostinelli et al., ). MCNP excels in handling fission trends over whole populations by using an averaged fission model (Verbeke et al., ). There are tools to simulate detector response with high fidelity within MCNP, such as MCNPX-Polimi (Clarke et al., ), but simulating multi-detector coincidence apparatus with charged particles remains a challenge. Geant4, developed by CERN, specializes in the simulation of tracking the transport of discrete simulated particles through matter, also using Monte Carlo methods. Utilizing object-oriented programming, the toolkit manages individual particles, tracking de-excitations and material interactions within defined geometries (J. Allison et al., ; Allison et al., ). Geant4 has superior ability to handle angular correlations, accurate nuclear de-excitations, and detector responses modeled on an event-by-event basis, enabling algorithmic development that can replicated experimental data analysis with high fidelity.
One notable weakness of Geant4 is the physics libraries for sampling fission fragments and trends in mass, energy, momenta, etc resulting from fission. Geant4 normally handles nuclei (and daughter particles) sequentially, while fission is inherently multi-nuclear. There are tools that can produce realistic samples of fission physics, but these tools have varying degrees of implementation in Geant4 (Hecht et al., ). Moreover, it is extremely computationally demanding to require Geant4 to simulate experiments involving large neutron-induced fission targets, particularly when accounting for incident neutrons. Following fission, the resultant fission fragments undergo complex chains of de-excitation and decay, while new fission events continue to occur. This leads to the overlapping of observables from both prior and newly initiated fission events at overlapping times, complicating data collection with incorrect false coincidences and increasing the challenge of accurately associating detected signals with their corresponding fission events. Simulating this process requires time-correlated data to effectively replicate overlapping false coincidences that must be managed in real-life experiments or repeated re-sampling of gamma spectra. Achieving such temporal correlations is difficult within Geant4’s event-by-event simulation framework, which is not inherently designed to handle the continuous time evolution of background radiation across multiple fission events.
There are suites of fission physics simulation tools that are not natively integrated into Geant4, such as GEF (Schmidt et al., ), TALYS (Fujio et al., ), and CGMF (Talou et al., ). These codes simulate known fission observables with high fidelity for discrete events. The authors suggest that the strengths of such tools can be paired with the strengths of Geant4, providing simulated detector responses and data analysis testing. The work herein utilizes CGMF, a computational fission physics tool published by LANL, that includes all the most relevant physics observables produced in fission (Talou et al.). JANGOFETT: A Novel Geant4-Operated Fission Event Tracking Tool has been developed to enable Geant4 simulations that incorporate covariant observables important for experimental work. This work presents a general overview of JANGOFETT, the implementation of CGMF outputs into Geant4, and a verification study conducted with 252Cf using JANGOFETT to indicate multiple overlapping events occur in the simulated ”experimental time”.
2 JANGOFETT Simulation Process
The process used in JANGOFETT is as follows. A CGMF simulation is initiated, generating fission fragments (pre-neutron evaporation) and primary fission products (”PFP”, post-neutron evaporation but before decaying) per the literature definitions (Madland, ). These are listed in a parsed csv file (which can be alternately used from a prior CGMF simulation for faster simulations). GEANT4 then iterates through the fission event list, and simulates each PFP on a particle-by-particle basis, with separate PFPs from the same fission event in successive simulations. After all simulations are completed, the detector hit times are time-shifted based on a rate sampling function (a randomly-sampled Poisson distribution, with the input of an assumed fission rate). Radiation from individual PFPs in the same fission event are time-shifted by the same amount to produce an ”experimental time” for correlations. The time-dependent detector responses are then checked for overlapping energy deposition steps within the same detector volume and time windows, then summed together to produce hits at the end of the Geant4 run. A rolling coincidence window is then implemented to associate detector hits with one another, and an analysis routine is employed. False coincidences occur when decays occur long after the prompt event that generated the PFPs. Further detail regarding this is described in a cartoon in Figure 1, as well as in the sections below.
2.1 Fission Simulation and Parsing
CGMF simulates fission fragment de-excitation and prompt decays immediately after nuclear scission, and readers are referred to the literature for details beyond the scope of JANGOFETT (Talou et al., )). CGMF is currently used to create the simulated fission fragments, as the list-based output of individual nuclei with properties provided before and after neutron evaporation makes it an ideal source of individual fission fragments for further processing. (At this time, the Hauser-Feschbach de-excitation radiation simulated are truncated before injection into Geant4, and the built-in Geant4 physics packages are used for de-excitation from a highly-excited nuclear state.)
Figure 1: Cartoon demonstrating how JANGOFETT receives and processes data. For each pair, each product is simulated as its own event.
The CGMF output is then parsed to provide information on the fragment after it has undergone neutron evaporation to Geant4. CGMF calculates much of this itself, and the PFP momentum vectors, kinetic energy, charge, mass, neutron multiplicity, and neutron energy are immediately processed into a new file containing the relevant information. However, the final excitation energy of the PFPs is also required, which is calculated in the parser as:
$\begin{array}[]{c}EX_{PFP}=KE_{I}+EX_{I}+[A(m_{amu})+\Delta_{FF}]\\ \quad-N(m_{N})-[(A-N)(m_{amu})+\Delta_{PFP}]-KE_{F}-\Sigma KE_{N}\end{array}$ (1)
Here, EX and KE refer to the excitation and kinetic energies, A is the PFP nuclear mass number, $m_{amu,N}$ are the mass of 1 amu and 1 neutron, $\Delta_{FF,PFP}$ is the mass defect for the ground state nucleus of the outgoing fission fragment (FF, pre-neutron evaporation particles) and PFP, N is the number of evaporated neutrons, and $KE_{N}$ represents the kinetic energies of the evaporated neutrons. The mass defects were retrieved from a repository (Huang et al., ). The emitted neutrons are not currently processed to be fed into Geant4, although their momentum transfer to the PFPs is incorporated, as our simulated experiment did not include neutron detection capabilities. This may be implemented in future versions. It should be noted that one nucleus pair (139Sb with one neutron evaporation from a FF of 140Sb and 112Ag) resulted in negative excitation energy by this method, with an average of -0.32 MeV. This error occurred 58 times in of $8.7\times 10^{7}$ fission events, but 100% of the time the PFP pair occurred, the error arose. The parser code forced the negative excitation energies to zero.
2.2 Geant4 Simulation
The PFP mass, momentum, and excitation energy list is passed to Geant4, which simulates each nucleus with input kinetic energy, excitation energy and momentum vectors. Geant4 physics processes incorporate statistical de-excitation models for internal transitions, allowing PFPs to be used as primary particles in the simulation.
The G4ParticleGun particle generator fires primary particles into the world volume at a given location, currently the center of the simulated detector array. Most often in Geant4, the particle gun is given a single particle (eg: a nucleus, nucleon, or a fundamental particle) and the desired number of simulations that the particle should be fired and tracked for. JANGOFETT uses the G4ParticleGun in such a way that, within a single run, different nuclei (read from the CGMF-generated and pre-parsed csv file) are fired for every event in the run. Pairs of particles are labeled with a fission event ID to facilitate time-shifting and debugging.
As each particle is fired, Geant4 begins a timeline of energy deposition events at time $t=0$ $(t_{0})$, with each following ‘step’ of the particle or its secondaries associated with a time referenced to this moment $t>t_{0}$. The fired PFP and subsequent radiation will proceed to interact with the world as defined by geometry until it leaves the world volume. After the particles interact with the world, they will continue to decay to stability, or until they reach a defined time limit. By default, the maximum particle lifetime of $10^{6}$ seconds has been set due to our specific interests in prompt emissions (and computational efficiency), but this setting could be changed to decay until stability if desired.
A text geometry file is expected for detector definitions, though modifications can be made for those familiar with Geant4. Simulations and analyses presented here use a detector apparatus with a pair of silicon charged-particle detectors in a mocked-up vacuum chamber surrounded by a thin polyimide film. The vacuum chamber is surrounded by high-purity germanium (HPGe) and bismuth germanate (BGO) veto detectors for $\gamma$-spectroscopy of varied geometries. This experimental apparatus is a useful simulation tool for correlating different radiation produced in fission and demonstrating that radiations are produced by Geant4 in a manner that enable multiple G4ParticleGun events to be associated.
2.3 Detector Hit Processing
In standard Geant4, the time stamp for particles created will always begin at $t_{0}$ in each event. This is not useful for simulations where false coincidences between fission events and their decays are expected at a meaningful rate. At the end of each event, the fission event ID of the particle is considered. Before Geant4 is called by the JANGOFETT script, a list of time shifts is created using a Poisson distribution based on the expected reaction rate and number of fission events in the simulation. These times shifts increase sequentially from an overall experimental time $t=0$ and represent the relative time of a fission event in the simulated experiment. This list is accessed by Geant4 and time shifts are applied to their fission event, creating an experimental time axis.
Once time shifted, the individual steps are summed into hits if they are within a modest time window (20 ns in the default implementation). Randomly sampled Gaussian blurs to both signal time and deposited energy are applied to the hits, based on the resolutions of the volume in which the energy deposition was recorded. Only the hits are saved, and the steps discarded, reducing file size by over 99%. At the end of the Geant4 run, the valid non-vetoed hits are analyzed together. Hits occurring within the coincidence time window of each other within the same detector volume are summed together, mimicking the response of detectors to multiple incident radiation. These finalized hits are then saved to a structured ROOT output file, which saves the energy deposition, detector volume, and the experimental time of the detection.
2.4 Novelty
While modifications have been made to Geant4 before to improve fission simulations, there has yet to be an implementation of a ”list-mode-like timeline” in which all fission events, from the beginning to the end of an experiment, are ordered chronologically. Previous modifications regarding fission have prioritized the correction of inaccuracies in the default fission models (Constantin et al., ) and improving the accuracy of isotopes, photons, and neutrons produced for use in statistical analysis of a large group of individual events (Tan and Bendahan, ). Though these are extremely useful themselves in certain contexts, JANGOFETT introduces the ability to replicate realistic detector outputs for fission experiments from external simulation tools.
2.5 Limitations
There are several additional parameters output by CGMF that are important in fission physics experiments that are not yet managed by JANGOFETT. Data such as the spin and parity are not currently parsed into Geant4. Other particles, such as the neutrons emitted and the gamma rays are not introduced into the simulated world volume, because the default configuration does not include neutron detection abilities, and Geant4 will automatically emit the gamma rays based on the excited state of the nuclei once fired. JANGOFETT can be applied to any fissile or spontaneous fissioning isotope, but the current implementation does not account for momentum transfer from high-energy neutrons (incident thermal neutrons $p_{n}\approx 0$), which influences the PFP input momentum vectors. This limitation restricts verification to spontaneous fission and thermal neutron induced fission cases. The charge and deformation of the fragments are also yet to be implemented. These missing parameters and limitations will be resolved in future updates to JANGOFETT.
Due to current computational limitations, the test run described in later sections experiences small numbers of events, resulting in low statistics. This will be easily managed with the upcoming Collaborative Innovation Complex to be built at Oregon State University. Much larger datasets may be simulated using the new supercomputer, which will provide data beyond what has been produced by this test system. The primary bottleneck is the simulation of fission in CGMF, which is responsible for over 90% of the computational time. More data are currently being simulated, improving statistics for further studies.
3 Verification
JANGOFETT has been benchmarked to ensure that the results it provides are physically meaningful when compared to evaluated data and consistent with input CGMF values. A test geometry was used comprising of 21 high purity germanium (HPGe) detectors with bismuth germanium oxide (BGO) active Compton veto. This is visualized and in supplementary material Figure S.1. Simulations presented here used a 252Cf spontaneous fission source, modeled with a fission rate of 1000/second and just over 24 hours of simulated experimental time. The JANGOFETT output that is used for verification contains only the timing, energy, and detector information, analyzed by a C++ CERN-ROOT code (Brun et al., 2019). This decision was motivated by the internal use-case for JANGOFETT, emulating the data structure from fission experiments, which output detector number, energy deposition, and time information for analysis.
3.1 Dataset
Figure 2: Time between coincident fission fragments pulses within silicon detectors. Note that a randomly-sampled Gaussian shift was applied to timing signals to represent uncertainties. If there is negligible time uncertainty this shape would shift to the right with more of a bell distribution.
(a) Kinetic Energy from coincident PFP energy depositions in silicon detectors.
(b) Total Kinetic Energy from sum of coincident PFP energy depositions in silicon detectors.
Figure 3: KE (a) and TKE (b) for the dataset with Gaussian fits in red. KE was determined by the energy in each silicon detector, considering fragments coincident within 20 ns. These KE were summed to produce TKE for coincident events.
Figure 4: Mass estimates based on KE from the 2E-method. The data are not unfolded or otherwise modified after the 2E loop is completed.
The test setup consisted of a system with two silicon detectors for measuring total kinetic energy (TKE) of fission fragments, surrounded by 21 HPGe detectors of varying geometries $\sim$20 cm from the source. Annular bismuth germanate (BGO) detectors surround the HPGEs for Compton suppression, which has been employed in the analysis used here.
Coincidence times between heavy fragments in silicon detectors are shown in Figure 2 and the coincidences between heavy fragments in the silicon detectors and gammas are shown within the supplemental material Figure S.2 , with a long tail that ends before 20 ns; beyond this point, false coincidences are about as likely as true. The average time for the two silicon detector signals was chosen as $\Delta_{time}=0$, and 50 ns coincidence windows (-30 ns to +20 ns) were chosen.
3.2 Kinetic Energy Simulations
Once coincidence windows were established, TKE was determined by summing the kinetic energies of coincident fission fragments. The resulting TKE distribution was fitted with a Gaussian, peaking at 180.083$\pm$0.001 MeV. Individual kinetic energy (KE) spectra were fitted with two Gaussians, with peaks at 79.544$\pm$0.002 and 102.229$\pm$0.001 MeV, which is within a reasonable calculational uncertainty of existing measurements for 252Cf as a calibration standard (181.03 MeV, 78.42 MEV, and 102.61) (Weissenberger et al., ). These are consistent with the CGMF outputs, indicating no bias in our analysis and steps-to-hits processing, other than being slightly low due to energy loss by X-rays and delta electrons.
Figure 3 shows a pair of KE (individual PFPs) and TKE (summed PFP energy) plots. Silicon detector signals with less than 20 MeV were excluded. As can be seen in Supplementary Figure S.3, there are a number of TKE signals that occur at $\sim$90 MeV and $\sim$180 MeV above the TKE peak - these are from false silicon detector triple and double coincidences that are present at a physically meaningful rate when considering the coincidence window. The TKE was analyzed before and after Geant4 simulations to rule out systematic errors -see the supplementary material for more detail (Figure S.4).
The fission fragment kinetic energy was further used to approximate fragment mass via the double-energy (2E) method (Schmitt et al., ), yielding mass distributions with an accuracy of approximately ±4 atomic mass units (amu). Figure 4 demonstrates the 2E-method mass yields without any additional gating conditions beyond timing coincidence.
3.3 252Cf Fragment Identification
With these coincidence criteria established, specific fission fragment pairs were selected for verification studies. Four heavy fragments were selected for gating, 144Ba, 141Cs, 149Nd, and 150Ce; the first is in the text of this article, while the remainder are included in the Supplementary Material. To isolate 144Ba (or 141Cs or 149149{149}149Nd or 150Ce) events, an initial mass gate was applied, selecting events within 4 amu of the desired nucleus, based on the 2E mass estimate. A second coincidence gate was then imposed, requiring $\gamma$-$\gamma$ coincidences from the heavy fragment to be measured in the HPGe detectors without a BGO Compton veto. $\gamma$-rays for coincidence were selected from a subset of high-probability $\gamma$-rays in the established nuclear data. Finally, an additional $\gamma$-gate was applied for each of the 0-5 neutron light fragment pairs. For these selected fragments, $\gamma$-$\gamma$-$\gamma$ spectroscopy histograms were generated to demonstrate both heavy and light fragments were simultaneously detected. It should be noted that gamma gating was not symmetric, e.g. if two gammas from the molybdenum isotope were identified but only one of the 144Ba, it was not included. Such a protocol would significantly increase the number of verified events, which was unnecessary to demonstrate the timing overlap of the PFPs, as was the goal here. Furthermore, no Doppler broadening was applied to energies for the purposes of analysis. Once applied, it is expected to improve statistics.
All peaks in the histograms were evaluated to determine whether the identified fragments could be plausibly misidentified and should be discarded. Our evaluations did not determine that any fragments needed to be discarded in this way, as all gamma signals that were well above background were able to be identified with reasonable certainty as coming from the heavy or light fragment. The results from the photon coincidence plots are shown in Figure 5 (see Figure S.5 - S.7 for other gated fragments and Table 1-4 for a list of the $\gamma$ energies used to gate in the supplemental material).
(a) 0n
(b) 1n
(c) 2n
(d) 3n
(e) 4n
(f) 5n
[…]
After identifying fragments from mass estimates and coincident $\gamma$, the KE of each fragment was summed to produced TKE plots for each of the light fragments observed, which are displayed in Figure 6.
(a) 0n
(b) 1n
(c) 2n
(d) 3n
(e) 4n
(f) 5n
[…]
4 Summary
JANGOFETT is a new tool for Geant4, developed to read primary fission product pairs into Geant4 with simulated time-based correlations. JANGOFETT performance was verified using a testbed of 252Cf(sf). Verification that the simulation produces distinguishable list-mode time sorted hits in detector volumes was performed by isolating coincidences from specific heavy and light fragments, based on mass estimates and photon energy. A heavy fragment of ${}^{1}44$Ba and its 0-5n light complements are shown here, with gamma-ray signals for each partner fragment appearing in the same simulation time, though they came from individual Geant4 instances. Additional fragments were analyzed and are included in the Supplementary Material. This tool enables the simulation and analysis of covariant fission observables in Geant4, generating data structures that are similar to those produced by experimental apparatus in fission for analysis.
JANGOFETT 1.0 has been made available at https://github.com/AlchemeyLab/JANGOFETT under GNU General Public License v3.0.
Acknowledgments
Authors LW and JS acknowledge fellowships from the Nuclear Regulatory Commission University Nuclear Leadership Program administered through Oregon State University. Authors gratefully acknowledge a startup package from the Oregon State University College of Engineering.
CRediT authorship contribution statement
L. Walker: Methodology, Software, validation, writing - original draft, writing - review and editing, visualization. J. Shire: Software, validation, writing - original draft, writing - review and editing, visualization. J. Jaffe: Software, validation, Writing - review and editing, visualization. P. Sprando: Software, validation, Writing - review and editing, visualization. J. Olinger: Writing - review and editing, visualization. A. Chemey: Conceptualization, methodology, software, writing - original draft, writing - review and editing, supervision, project administration.